OpenStudy (mathisfun13):

find the geometric series with r=3 and s=364

OpenStudy (dumbcow):

\[s_n = \frac{a_1(r^{n} -1)}{r-1}\] \[364 = \frac{a_1(3^{n} -1)}{2}\] need more info to find a_1 and n

OpenStudy (mathisfun13):

thats all they give u

ganeshie8 (ganeshie8):

you may do trial and error then

ganeshie8 (ganeshie8):

1 + 3 + 3^2 + ....

OpenStudy (mathisfun13):

oh the question says find a1 for the geometric series

OpenStudy (dumbcow):

you need a reference point, at least one number in the sequence

ganeshie8 (ganeshie8):

see when the sum reaches 364

OpenStudy (mathisfun13):

how so

ganeshie8 (ganeshie8):

since there is not enough info, this is not a direct q... they may want you test few sums and figure out the series

ganeshie8 (ganeshie8):

if psble, take a screenshot of the q u seeing and attach... it helps to confirm :)

OpenStudy (mathisfun13):

mmkaay

ganeshie8 (ganeshie8):

Now that makes sense. Did you notice, you missed the 6 in \(S_6\)

OpenStudy (mathisfun13):

lol ohhh no i didnt sorryy

ganeshie8 (ganeshie8):

its okay :)

ganeshie8 (ganeshie8):

So, you're given r=3, and n = 6.

ganeshie8 (ganeshie8):

plug them in the formula given by dumbcow and solve \(a_1\)

OpenStudy (mathisfun13):

okay

ganeshie8 (ganeshie8):

\(\large s_n = \frac{a_1(r^{n} -1)}{r-1}\) \(\large s_6 = \frac{a_1(3^{6} -1)}{3-1}\) \(\large 364 = \frac{a_1(3^{6} -1)}{2}\)

ganeshie8 (ganeshie8):

you can solve \(a_1\)

OpenStudy (mathisfun13):

would it be 243?

ganeshie8 (ganeshie8):

it wont be that big for sure

ganeshie8 (ganeshie8):

try again, use ur calculator

OpenStudy (dumbcow):

oh you are given n ... i was going to say there are multiple answers given an "n" you can find a "a_1" that works

OpenStudy (mathisfun13):

idk i got 364

OpenStudy (mathisfun13):

the a1 is just there?

ganeshie8 (ganeshie8):

\(\large 364 = \frac{a_1(3^{6} -1)}{2} \)

ganeshie8 (ganeshie8):

multiply 2 both sides

ganeshie8 (ganeshie8):

\(\large 364 \times 2 = a_1(3^{6} -1) \)

ganeshie8 (ganeshie8):

\(\large \frac{364 \times 2}{(3^6-1)} = a_1\)

ganeshie8 (ganeshie8):

wolfram says it equals 1 http://www.wolframalpha.com/input/?i=%28364*2%29%2F%283%5E6-1%29